I am so confused with how to figure out the formula to find the answer when it asks me to graph x-y=3. I just don't know how or where to begin. Do I make up a number to use?
Answer by Fombitz(32388)

x-y=3
y=x-3
So now you have a formula for the dependent variable y given the independent variable x.
It's the equation of a line so choose two values of x and get two data points.
Then graph.
When x=0, y=0-3=-3
(0,-3)
When x=3, y=3-3=0
(3,0)
Graph it.
.